OpenStudy (anonymous):

Prove: 1/cosec x -cot x = cosec x +cot x

OpenStudy (anonymous):

stuck here:-

OpenStudy (anonymous):

sin x /1-cos x

hartnn (hartnn):

no...in first step , just multiply and divide by cosec x+cot x

hartnn (hartnn):

then use, cosec^2x-cot^2 x=1 standard identity

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

cos^4x - sin^4x what will be starting of this?? @hartnn

hartnn (hartnn):

(c^2+s^2)(c^2-s^2) = 1*(c^2-s^2)

hartnn (hartnn):

= cos 2x

OpenStudy (anonymous):

cos^4x -sin^4x =2cos^x-1

OpenStudy (anonymous):

I am here :- (cos^2x-sin^2x)

OpenStudy (anonymous):

(cos x + sin x)(cosx - sin x )

OpenStudy (anonymous):

is this correct?

hartnn (hartnn):

what u want to prove it to be ?

OpenStudy (anonymous):

2cos^2x -1

hartnn (hartnn):

cos^2 -sin^2 use, sin^2 =1-cos^2

hartnn (hartnn):

because sin^2 +cos^2 =1

OpenStudy (anonymous):

thanks!!

hartnn (hartnn):

welcome :)